Ticket Number: HEA-NOC/20100526-5 Ticket Status: UPDATE
Ticket Type: scheduled Resolver: HEA-NOC
Ticket Opened: 20100526 13:59 UTC+1 Problem Start: 20100527 07:30 UTC+1
Ticket Update: 20100624 09:06 UTC+1 Problem End: 20100527 09:00 UTC+1
TCD are being upgraded 10Gig links.
The cutover from teir existing 1Gig links to the 10Gig links is to take
Thursday 27th May at 07:30. A window of 1.5 hours is scheduled.
20100526 13:57 GL There will be a brief outage as TCD need to do some
repatching wtihin their LAN and for BGP convergence.
20100527 08:44 GL TCD migrated to 10Gig links at 07:35. TCD internal
testing concluded at 08:40. TCD 10Gig links in full production. Monitoring
20100615 10:33 GL Failover testing to occur on Thursday 17th June at 08:00
20100615 10:34 GL Intermittent outages to TCD as failover as testing take
20100617 08:36 GL Failover testing completed successfully. Next stage is
to remove previous 1Gb links and cleanup monitoring.
20100624 09:05 GL Removed old 1Gig links from configuration and
monitoring. Next step is to officially sign off on the project
This ticket can be monitored at http://www.hea.net/tickets/20100526-5
HEAnet Limited. Registered in Ireland, No. 275301.